# Training a multilayer dynamical spintronic network with standard machine learning tools to perform time series classification

### v2

@@ VI Conclusion
We have simulated a multi-layer neural network where the neurons are spintronic oscillators. Leveraging the transient dynamics of the neurons for memory and non-linearity, we have performed a time series classification task, sequential digits. We demonstrate that this multi-layer dynamical spintronic network can be trained with backpropagation through time (BPTT) and standard machine learning tools, which is key for scaling to more complex tasks. Despite having a bounded internal variable, the spintronic network achieves $89.83~{}\%$ accuracy, as good as a standard recurrent neural network. These results demonstrate the potential of spintronic oscillators for time-series processing. We have devised guidelines for the design of such networks and their training: the relaxation time must be longer than the time-scale of the input and the cumulative drive must be close to one. Once built, a dynamical spintronic network can be trained to learn tasks over a 5-fold range of input time-scales. We have shown that we can sparsify the network down to $50~{}\%$ connectivity before training, further reducing the footprint and energy consumption of the system. For the sequential digits tasks and the considered architecture, we estimate that the network would consume about 40 pJ per image. These results open the path to the scaling up of spintronic neural networks for time-series processing at low energy cost.

### v3

@@ VI Conclusion
We have simulated a multi-layer neural network where the neurons are spintronic oscillators. Leveraging the transient dynamics of the neurons for memory and non-linearity, we have performed a time series classification task, sequential digits. We demonstrate that this multi-layer dynamical spintronic network can be trained with BPTT and standard machine learning tools, which is key for scaling to more complex tasks. Despite having a bounded internal variable, the spintronic network achieves $89.83~{}\%$ accuracy, as good as a standard recurrent neural network. These results demonstrate the potential of spintronic oscillators for time-series processing. We have devised guidelines for the design of such networks and their training: the relaxation time must be longer than the time-scale of the input and the cumulative drive must be close to one. Once built, a dynamical spintronic network can be trained to learn tasks over a 5-fold range of input time-scales. We have shown that we can sparsify the network down to $50~{}\%$ connectivity before training, further reducing the footprint and energy consumption of the system. For the sequential digits tasks and the considered architecture, we estimate that the network would consume about 40 pJ per image. Despite the large interest for time-series processing with physical oscillators – whether spintronic, photonic, resistive or CMOS-based – there is little work on systematic scalable training methods for oscillators. Here, we show how to use standard machine learning tools to train multilayer spintronic oscillators, a method that is scalable. We also point on important design rules regarding the matching of the network dynamics with the input characteristics. Our use of a simplified model that is general enough makes our work easily adaptable to other types of physical oscillators. In this work we have demonstrated the concept with a simple task. Future studies will tackle larger networks and state of the art machine learning tasks. These results open the path to the scaling up of spintronic neural networks for time-series processing at low energy cost.
@@ Acknowledgements
Appendix
We study the impact of the connection density on the overfitting. Figure 5 shows the evolution of the accuracy and loss versus the number of epochs, for the train and test datasets at different global connection densities. The conditions are the same as described in Section III. We observe that the gap between the train and test datasets decreases with the connection density. This shows that reducing the density helps to prevent overfitting and explains why the optimal density is 50% rather than 100%.
Figure 5: Right: Evolution of the accuracy (in % of correctly classified inputs) versus the number of epochs. Left: Evolution of the loss versus the number of epochs. The train dataset is represented with solid lines and the test dataset with dashed lines. The colors blue, green and red correspond to global connection densities of 50%, 25% and 10% respectively.